OpenStudy (anonymous):

show that lim x approaching +infinity ([root ([x sqrt]+x )+x)=1/2

OpenStudy (experimentx):

?? can't understand that function

OpenStudy (anonymous):

(sqrt root (xsquared +x) +x = 1/2

OpenStudy (anonymous):

square root

OpenStudy (experimentx):

\[ \sqrt{x^2 + x} + x\] ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

does that not just diverge? unless i misunderstand

OpenStudy (anonymous):

i know i am supposed to rationalize

OpenStudy (experimentx):

wolf bro says no http://www.wolframalpha.com/input/?i=lim+x- >inf+sqrt%28x%5E2+%2Bx%29+%2B+

OpenStudy (turingtest):

might as well try to multiply by the conjugate

OpenStudy (anonymous):

yes but after rationalizing dont what to do

OpenStudy (experimentx):

There might be some error with the question, the limit clearly diverges to infinity

OpenStudy (experimentx):

*I think

OpenStudy (anonymous):

- infinity

OpenStudy (anonymous):

ok thxs

OpenStudy (turingtest):

yeah, this has to diverge... but...\[\sqrt{x^2+x}+x\cdot{\sqrt{x^2+x}-x\over\sqrt{x^2+x}-x}= {x^2+x-x^2\over\sqrt{x^2+x}-x}={x\over\sqrt{x^2+x}-x}\]\[{1\over\sqrt{1+\frac1x}-1}\]taking the limit gives\[\frac10\]so we are going to positive infinity it looks like to me...

OpenStudy (experimentx):

most likely, your function is http://www.wolframalpha.com/input/?i=lim+x-%3E+infinity++sqrt%28x%5E2%2Bx%29+-+x

OpenStudy (turingtest):

^that makes sense

OpenStudy (experimentx):

In that case, Follow Turing Test, instead of - on that left, put + ... everything is almost same.

OpenStudy (experimentx):

Thanks everyone .. for making Guru!!

OpenStudy (turingtest):

welcome :) well-deserved!

OpenStudy (anonymous):

i agree, well earned. you can solve the actual problem using the same method @TuringTest used on the other function

OpenStudy (anonymous):

What is the correct statement?

OpenStudy (anonymous):

it doesnt work. it just goes to positive infinity

OpenStudy (freckles):

\[\frac{\sqrt{x^2+x}+x}{1} \cdot \frac{\sqrt{x^2+x}-x}{\sqrt{x^2+x}-x}=\frac{(x^2+x)-x}{\sqrt{x^2+x}-x}\] \[=\frac{x}{\sqrt{x^2+x}-x} \cdot \frac{\frac{1}{\sqrt{x^2}}}{\frac{1}{\sqrt{x^2}}}= \frac{\frac{x}{\sqrt{x^2}}}{\frac{\sqrt{x^2+x}}{\sqrt{x^2}}-\frac{x}{\sqrt{x^2}}}\] |x|=x since x>0 Recall \[\sqrt{x^2}=|x|\] Ok then so we have \[\frac{\frac{x}{x}}{\sqrt{1+\frac{1}{x}}-\frac{x}{x}}\] \[\lim_{x \rightarrow \infty}\frac{1}{\sqrt{1+\frac{1}{x}}-1}=\frac{1}{\sqrt{1+0}-1}=\frac{1}{1-1} \text{ DNE} \] Sorry had to do it myself lol

OpenStudy (turingtest):

that would be positive infinity, not DNE

OpenStudy (freckles):

so infinity is a number ?

OpenStudy (freckles):

infinity does not exist

OpenStudy (turingtest):

no, but the limit not existing is different than one that goes to infinity

OpenStudy (turingtest):

if it goes to -infty from the left and +infty from the right, then the limit DNE

OpenStudy (freckles):

You can say either one

OpenStudy (turingtest):

no you cannot, I assure you

OpenStudy (freckles):

Yes you can All numbers exist infinity is not a number so it does not exist

OpenStudy (turingtest):

typo, meant this one dne http://www.wolframalpha.com/input/?i=lim+x-%3E+0++x%2F%7Cx%7C

OpenStudy (turingtest):

\[\lim_{x\to\infty}x=+\infty\]\[\lim_{x\to\infty}-x=-\infty\]\[\lim_{x\to0}{x\over|x|}\text{DNE}\]

OpenStudy (turingtest):

all different

OpenStudy (turingtest):

only if\[\lim_{x\to a^+}f(x)\neq\lim_{x\to a^-}f(x)\]does the limit not exist

OpenStudy (anonymous):

give some explain here @TuringTest http://openstudy.com/study?source=email#/updates/4f8b04ede4b09e61bffc5631

OpenStudy (freckles):

If both sides go to positive infinity you can say the limit is positive infinity but you can also say the limit does not exist because infinity is not a number Same thing for if both sides go to negative infinity Sure the positive infinity/negative infinity gives us more information about the behavior of the graph but we can say the limit does not exist because infinity is like totally not a number And of course we can say the limit does exist when the left limit doesn't equal right limit But we can not say the limit is infinity if the left limit doesn't equal the right limit. So in all cases where you have the limit is infinity (+/-) you can say DNE But not in all cases where you have the limit does not exist can you say the limit is infinity.

OpenStudy (turingtest):

infinity is not a number, but if you say that that means\[-\infty=\infty=DNE\]you are saying that all three functions I wrote above have the same limit try taking that to your professor and see what they tell you I think you are confusing divergence with non-existence for limits

OpenStudy (freckles):

No that is not what I'm saying

OpenStudy (freckles):

I didn't say -inf=inf

OpenStudy (freckles):

I said if you have -inf then you can say the limit does not exist because -inf is not a number I said if you have inf then you can say the limit does not exist because inf is not a number I never ever said -inf=inf

OpenStudy (turingtest):

I think we should post this as a separate question and get more opinions "If the limit of a function is +/-infinity, does that mean that the limit DNE" lets ask!

OpenStudy (anonymous):

this is the definition i know, if we're going by definitions The limit of f(x) as x approaches a is L if and only if, given e > 0, there exists d > 0 such that 0 < |x - a| < d implies that |f(x) - L| < e